# Spectral Moment Formulae for $\hbox{GL}(3)\times\hbox{GL}(2)$ L-functions II: The Eisenstein Case

### v2

@@ Abstract.
This work is the second in a series, following Part I [Kw23] (Algebra Number Theory 18.10 (2024)) and preceding Part III [Kw25] (Math. Ann. 391.1 (2025)). We continue our investigation of spectral moments of $\hbox{GL}(3)\times\hbox{GL}(2)$ $L$-functions from the perspective of period integrals. Using an identity between two distinct periods for the $\hbox{GL}(3)$ Eisenstein series, we establish an exact Motohashi-type identity linking the shifted cubic moment of $\hbox{GL}(2)$ $L$-functions to the shifted fourth moment of $\hbox{GL}(1)$ $L$-functions. In addition, we offer a novel, intrinsic and automorphic account for the sources and symmetries of the full set of main terms for both moments, in agreement with the CFKRS Moment Conjectures (Proc. Lond. Math. Soc. (3) 91 (2005)).
@@
Automorphic forms, Automorphic $L$-functions, Rankin–Selberg $L$-functions, Riemann zeta-function, Period integrals, Maass forms, Eisenstein series, Poincaré series, Moments of $L$-functions, CFKRS conjectures, Spectral reciprocity, Motohashi’s formula, Whittaker functions, Hypergeometric functions, Barnes integrals
@@ 1.1. Background
Over the past two decades, the cubic moment of $\hbox{GL}(2)$ automorphic $L$-functions (or simply the cubic moment) has driven a wave of remarkable applications in number theory. Notable among these is the landmark result of Conrey–Iwaniec [CI00], along with the vast recent developments of Petrow–Young [PY19, PY20, PY23], culminating in the Weyl-strength subconvexity bounds for Dirichlet $L$-functions of any primitive character $\chi\,(\bmod\ q)$:
@@
and analogous bounds for $\mathrm{GL}(2)$ $L$-functions in various aspects. These advances have far-reaching arithmetic consequences; see, for instance, [Y17, DIT16, FM10, LMY13].
Structurally, it is striking that the cubic moment of $\hbox{GL}(2)$ $L$-functions is dual to the fourth moment of the Riemann zeta function (or the fourth moment)—a fundamentally different moment that occupies a central place in classical analytic number theory, tracing back to the seminal works of Hardy–Littlewood [HL16] and Ingham [Ing27] in the 1920s. This duality was first discovered by Motohashi [Mo93] in the 1990s, who established a deep spectral identity of the form
@@
where $w\mapsto\widetilde{w}$ denotes an intricate integral transform, and ()(***) encodes a collection of non-trivial arithmetic contributions together with the continuous component. This identity is widely regarded as a “crowning achievement” in the theory of the Riemann zeta function ([CK15a, CK15b]), and it represents the first and a particularly intriguing (and challenging) instance of Spectral Reciprocity.
To study the cubic moment, one proceeds by inverting the transform $w\mapsto\widetilde{w}$ and the identity (1.2). This was achieved to some extent (i.e., without revealing the underlying fourth moment) by Ivić [Iv01], and Motohashi sketched a formal derivation in his note [Mot99]. Proving a spectrally inverted Motohashi-type formula, even in approximate form, requires substantial refinements of the techniques of [CI00] and [Iv01]. Such refinements were obtained only relatively recently by Petrow [Pe15] and Frolenkov [Fr20], both via relative trace formulae. The relevant analytic machinery has been developed in various ways, e.g., [Bl12b, Li11, Y17, JM05, Qi24, KPY19].
@@ 1.2. Moment conjectures
A central theme in moments of $L$-functions is to understand the sources and structures underlying the full set of main terms. Over the years, this pursuit has drawn sustained attention and yielded fruitful outcomes. Most notably, two sets of heuristics have led to remarkably precise conjectures for the asymptotics of moments. Conrey–Farmer–Keating–Rubinstein–Snaith [CFK+05] (hereafter referred to as CFKRS) developed their “recipes” based on shifted moments and approximate functional equations, whose use dates back to [Ing27] and [HL16]. Their approach reveals surprising connections to Random Matrix Theory and brings to light many unexpected intrinsic symmetries and combinatorial structures. On the other hand, Diaconu–Goldfeld–Hoffstein [DGH03] introduced the Multiple Dirichlet Series for moments of the Riemann zeta function and quadratic Dirichlet $L$-functions. Their predictions arise from the (conjectural) analytic continuation, polar divisors, and the group of functional equations of these series.
For the two moments considered in this article, namely those in (1.2), extracting the full structure of the main terms, suppressed as part of ()(***) in (1.2), is a delicate task. For the fourth moment of the $\zeta$-function, Heath-Brown [HB79] and Motohashi [Mo93] proved the asymptotic formula
@@
with $\delta<1/8$ and $\delta<1/3$ respectively, where $P_{2}$ is a polynomial of degree $4$. The resulting expressions of $P_{2}$ (see also Conrey [C96]) were, unfortunately, quite complicated, making a complete determination of its coefficients daunting from their formulae.
In [Mo93], Motohashi considers the following smoothed and shifted variant of the fourth moment:
@@
The shifts are initially taken with $\operatorname{Re}\alpha_{i},\,\operatorname{Re}\beta_{j}\gg 1$, and play a pivotal role in an intricate analytic continuation procedure. CFKRS later discovered that the many degenerate terms in Motohashi’s formula for (1.4) can be combined, in rather subtle ways, to yield an elegant, intrinsic description of the main terms. There are six main terms in the CFKRS formulation (see (5.2)), from which the polynomial $P_{2}$ can be computed in full and with ease by letting the shifts $\alpha_{i},\beta_{j}\to 0$. This is explained in [CFK+05, Section 5.1].
The recipe works equally well for the cubic moment of $\hbox{GL}(2)$ $L$-functions. Owing to differences in the families and their arithmetic, this case yields eight distinct main terms. In [CI00], the authors located only the diagonal term, though they noted the interest in identifying the others ([CI00, p. 1177]). We review these conjectures in Section 5, and examine the difficulties that hindered the identification of these terms in earlier approaches in Section 1.4.
@@ 1.3. Beyond the Petersson/Kuznetsov paradigm
Motohashi’s original work [Mo93] is technical, and the source of “reciprocity” is far from apparent. A priori, there is no reason to expect the off-diagonal (or error) terms to be neatly expressible by moments of $L$-functions. Having a conceptual understanding of this phenomenon, without intermediate devices like Kloosterman sums, summation formulae, shifted convolutions, or the Petersson–Kuznetsov formulae, is therefore of significant interest.
This direction was perhaps first pursued by Bruggeman–Motohashi [BM05], whose approach to constructing the two distinct-looking moments in (1.2) involves a combination of spectral projections, Hecke operators, and a two-fold limiting process applied to an automorphic kernel. The choice of the kernel is by no means straightforward: it entails delicate smoothing and regularization, along with careful control of convergence, regularity, and interchange of limits. These ingredients, in turn, illuminate the integral transform through the theories of Whittaker–Kirillov models and Bessel functions for the group $\hbox{PGL}_{2}(\mathbb{R})$ (see [CPS90]).
Michel–Venkatesh [MV06, Section 4], [MV10, Section 4.5], and Reznikov [R08, p. 453] took a further step towards an account more intrinsic than [BM05]. They envisioned proving Motohashi’s formula within the structural framework of the strong Gelfand formation and via a “regularized” geodesic period for the product of two $\hbox{PGL}(2)$ Eisenstein series, formally given by
@@
which morally suggests a matching of the spectra $L^{2}([\mathrm{PGL_{2}}])$ and $L^{2}([\mathrm{GL_{1}}])$, resulting in Motohashi’s formula. However, numerous technical obstacles must be overcome to implement their strategy. In particular, the sources of various main (or degenerate) terms are not evident, and many integrals suffer from serious divergence issues due to the non-compactness of domains and the non-square-integrability of Eisenstein series. More refined regularization techniques than those developed in [MV10] are therefore necessary; see also [BHKM20] for other intricacies of this approach. Only very recently did Nelson [Ne20+] establish full rigour for the Michel–Venkatesh–Reznikov period integral strategy, through a range of highly impressive and sophisticated techniques.
Wu [Wu22] gave another interpretation of Motohashi’s formula by generalizing [BM05] with a pre-trace formula on the Schwartz space of 22-by-22 matrices, which incorporates the Godement–Jacquet and Tate $L$-functions on the spectral and geometric sides of the formula. Further explications and analytic subtleties were addressed in [BFW21+], which draws extensively on deep results in the theory of special functions. The authors of [BFW21+] discuss the obstacles to obtaining an exact spectral inversion of Motohashi’s formula.
@@ 1.4. Our main results
In this article, we present a new proof of the spectral inversion of Motohashi’s identity (1.2) via the $\hbox{GL}(3)$ period integral method of [Kw23]. Our approach offers fresh perspectives on the reciprocity phenomena and surmounts technical obstacles in earlier treatments. Furthermore, we elucidate the links between periods and the Moment Conjectures of [CFK+05].
In our previous work [Kw23] (resp. [Kw25]), we examine the (twisted) spectral first moment of $\hbox{GL}(3)\times\hbox{GL}(2)$ Rankin–Selberg $L$-functions with the $\hbox{GL}(3)$ automorphic form being a fixed Hecke–Maass cusp form. We uncover a higher-rank Motohashi-type identity that underpins Li’s celebrated convexity-breaking results [Li11].
This line of works belong to the recent framework Period Reciprocity, for which only a handful of other instances are known: [MV10, Bl12a, Ne20+, Nu20+, Za20+, Za21, JN21+, Mi21+]. The tools involved lie largely outside the traditional realm of analytic number theory, and they streamline the intricate chains of analytic-arithmetic transformations prevalent in the literature. We refer the reader to [Kw23, Sections 3-4], [Kw25, Section 1] and [JN21+] for more in-depth discussions of the implementations and technical features of this framework.
To state our main results, let $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ denote the following shifted cubic moment:
@@
where $(\phi_{j})_{j=1}^{\infty}$} is an orthogonal basis of even Hecke–Maass cusp forms of $\mathrm{SL}_{2}(\mathbb{Z})$ with respect to the Petersson inner product $\langle\,\,\cdot\,$$\cdot\,\rangle$, and satisfies $\Delta\phi_{j}=(1/4-\mu_{j}^{2})\phi_{j}$} (with $\Delta:=-y^{2}(\partial_{x}^{2}+\partial_{y}^{2})$). Let $W_{3}$ be the Weyl group of $\hbox{GL}(3)$, which consists of the permutation matrices
Let $\mathfrak{a}_{\mathbb{C}}:=\{(\alpha_{1},\alpha_{2},\alpha_{3})\in\mathbb{C}^{3}:\alpha_{1}+\alpha_{2}+\alpha_{3}=0\}$. The actions of $w\in W_{3}$ on $\mathfrak{a}_{\mathbb{C}}$ are given by:
To simplify our residual calculations (as in [CFK+05]), we introduce:
@@
In practical applications of any spectral summation formula, it is essential to have an ample supply of admissible test functions. The following regularity assumption serves our purposes and is satisfied by those commonly used in the literature (cf. [Kw23, Remark 5.27]).
@@
Let $h:(0,\infty)\rightarrow\mathbb{C}$ denote the inverse Kontorovich–Lebedev transform of $H$ (see Definition 3.8), $d^{W}\mu:=|\Gamma(\mu)|^{-2}\,(d\mu/4\pi i)$, $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\,\Gamma(s/2)$,  $d^{*}\mathbf{y}=(y_{0}y_{1}^{2})^{-1}\,dy_{0}\,dy_{1}$, $e(x):=e^{2\pi ix}$, $\Lambda(s)$ and $\Lambda(s,\,\phi)$ be the completed $\zeta$-function and $L$-function of a cusp form $\phi$, and $W_{-\alpha}(X,y_{1})$ be the standard (Jacquet) Whittaker function for $\mathrm{PGL}_{3}(\mathbb{R})$ (see Lemma 3.4).
@@ 1.4.1. The reciprocity formula
$\mathfrak{M}_{\alpha}^{(3)}(s;H)\ +\ \mathcal{D}_{\alpha}(s;H)\ +\ \mathcal{R}_{\alpha}(s;H)$
$\hskip 50.0pt\ =\ \sum_{i=1}^{3}\,\mathcal{M}_{\alpha}^{i}(s;H)\,+\ \frac{1}{2}\ \int_{(1/2)}\zeta\left(2s-s_{0}\right)\ \prod_{i=1}^{3}\ \zeta(s_{0}+\alpha_{i})\left(\mathcal{F}_{\alpha}H\right)\left(s_{0},s\right)\ \frac{ds_{0}}{2\pi i},$ (1.9)
@@
$(\mathcal{F}_{\alpha}H)(s_{0},s)\ :=\ \sum_{\pm}\,\int_{0}^{\infty}\int_{0}^{\infty}\,h\Big(\frac{y_{1}}{\sqrt{1+y_{0}^{2}}}\Big)\int_{0}^{\infty}W_{-\alpha}(X,y_{1})e(\pm Xy_{0})X^{s_{0}-1}\,d^{\times}X\,\frac{y_{0}^{2s-s_{0}}y_{1}^{s-\frac{1}{2}}}{(1+y_{0}^{2})^{\frac{s}{2}-s_{0}+\frac{1}{4}}}\,d^{*}\mathbf{y},$ (1.10)
$\mathcal{M}_{\alpha}^{0}(s;H)\,$ $:=\,\prod_{1\leq i<j\leq 3}\,\zeta(2s-\alpha_{i}-\alpha_{j})\ \int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i})\ d^{W}\mu,$ (1.11)
$\mathcal{M}_{\alpha}^{1}(s;H)\,$ $:=\,\sum_{w=w_{2},w_{4},w_{\ell}}\,\ \zeta(1+\alpha_{2}^{w}-\alpha_{3}^{w})\zeta(1+\alpha_{1}^{w}-\alpha_{3}^{w})\zeta(2s+\alpha_{1}^{w}+\alpha_{2}^{w})$
$\hskip 50.0pt\times\ \int_{(0)}\,H(\mu)\,\prod_{\pm}\ \Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{1}^{w})\Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{2}^{w})\Gamma_{\mathbb{R}}(1-s\pm\mu-\alpha_{3}^{w})\ d^{W}\mu,$ (1.12)
$\mathcal{M}_{\alpha}^{2}(s;H)\,$ $:=\,\ \sum_{i=1}^{3}\ \zeta\big(2-2s+\sum_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\alpha_{j}\big)\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\zeta(1-\alpha_{i}+\alpha_{j})$
$\hskip 50.0pt\,\times\,\int_{(0)}\,H(\mu)\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i})\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{j})\ d^{W}\mu,$ (1.13)
$\mathcal{M}_{\alpha}^{3}(s;H)\,$ $:=\,\ \prod_{1\leq i<j\leq 3}\,\zeta(2-2s+\alpha_{i}+\alpha_{j})\int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\ \prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{i})\,d^{W}\mu,$ (1.14)
$\mathcal{D}_{\alpha}(s;H)\ :=\ 2\sum_{w=w_{3},w_{5},w_{\ell}}\$$H(s-\alpha_{1}^{w})\ \frac{\Lambda(1-\alpha_{1}^{w}+\alpha_{2}^{w})\Lambda(1-\alpha_{1}^{w}+\alpha_{3}^{w})\Lambda(2s-\alpha_{1}^{w}-\alpha_{2}^{w})\Lambda(2s-\alpha_{1}^{w}-\alpha_{3}^{w})}{\Lambda(1+2(s-\alpha_{1}^{w}))},$ (1.15)
$\mathcal{R}_{\alpha}(s;H)\$:= $:=\ 2\ \sum_{i=1}^{3}\ \ H(1-s+\alpha_{i})\ \frac{\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\ \Lambda(1+\alpha_{i}-\alpha_{j})\Lambda(2s-1-\alpha_{i}-\alpha_{j})}{\Lambda\left(3-2s+2\alpha_{i}\right)}.$ (1.16)

### v3

@@ Abstract.
This work is the second in a series, following Part I [49] (Algebra Number Theory 18.10 (2024)) and preceding Part III [50] (Math. Ann. 391.1 (2025)). We continue our investigation of spectral moments of $\hbox{GL}(3)\times\hbox{GL}(2)$ $L$-functions from the perspective of period integrals. Using an identity between two distinct periods for the $\hbox{GL}(3)$ Eisenstein series, we establish an exact Motohashi-type identity linking the shifted cubic moment of $\hbox{GL}(2)$ $L$-functions to the shifted fourth moment of $\hbox{GL}(1)$ $L$-functions. In addition, we offer a novel, intrinsic and automorphic account for the sources and symmetries of the full set of main terms for both moments, in agreement with the CFKRS Moment Conjectures (Proc. Lond. Math. Soc. (3) 91 (2005)).
@@
Automorphic forms, Automorphic $L$-functions, Rankin–Selberg convolution, Riemann zeta-function, Period integrals, Maass forms, Eisenstein series, Poincaré series, Moments of $L$-functions, CFKRS conjectures, Spectral reciprocity, Motohashi’s formula, Whittaker functions, Hypergeometric functions, Barnes integrals
@@ 1.1. Background
Over the past two decades, the cubic moment of $\hbox{GL}(2)$ automorphic $L$-functions (or simply the cubic moment) has driven a wave of remarkable applications in number theory. Notable among these is the landmark result of Conrey–Iwaniec [24], along with the vast recent developments of Petrow–Young [65, 66, 67], culminating in the Weyl-strength subconvexity bounds for Dirichlet $L$-functions of any primitive character $\chi\,(\bmod\ q)$:
@@
and analogous bounds for $\mathrm{GL}(2)$ $L$-functions in various aspects. These advances have far-reaching arithmetic consequences; see, for instance, [74, 28, 29, 54].
Structurally, it is striking that the cubic moment of $\hbox{GL}(2)$ $L$-functions is dual to the fourth moment of the Riemann zeta function (or the fourth moment)—a fundamentally different moment that occupies a central place in classical analytic number theory, tracing back to the seminal works of Hardy–Littlewood [33] and Ingham [39] in the 1920s. This duality was first discovered by Motohashi [59] in the 1990s, who established a deep spectral identity of the form
@@
where $w\mapsto\widetilde{w}$ denotes an intricate integral transform, and ()(***) encodes a collection of non-trivial arithmetic contributions together with the continuous component. This identity is widely regarded as a “crowning achievement” in the theory of the Riemann zeta function ([19, 20]), and it represents the first and a particularly intriguing (and challenging) instance of Spectral Reciprocity.
To study the cubic moment, one proceeds by inverting the transform $w\mapsto\widetilde{w}$ and the identity (1.2). This was achieved to some extent (i.e., without revealing the underlying fourth moment) by Ivić [42], and Motohashi sketched a formal derivation in his note [61]. Proving a spectrally inverted Motohashi-type formula, even in approximate form, requires substantial refinements of the techniques of [24] and [42]. Such refinements were obtained only relatively recently by Petrow [64] and Frolenkov [30], both via relative trace formulae. The relevant analytic machinery has been developed in various ways, e.g., [10, 52, 74, 46, 68, 47].
@@ 1.2. Moment conjectures
A central theme in moments of $L$-functions is to understand the sources and structures underlying the full set of main terms. Over the years, this pursuit has drawn sustained attention and yielded fruitful outcomes. Most notably, two sets of heuristics have led to remarkably precise conjectures for the asymptotics of moments. Conrey–Farmer–Keating–Rubinstein–Snaith [22] (hereafter referred to as CFKRS) developed their “recipes” based on shifted moments and approximate functional equations, whose use dates back to [39] and [33]. Their approach reveals surprising connections to Random Matrix Theory and brings to light many unexpected intrinsic symmetries and combinatorial structures. On the other hand, Diaconu–Goldfeld–Hoffstein [26] introduced the Multiple Dirichlet Series for moments of the Riemann zeta function and quadratic Dirichlet $L$-functions. Their predictions arise from the (conjectural) analytic continuation, polar divisors, and the group of functional equations of these series.
For the two moments considered in this article, namely those in (1.2), extracting the full structure of the main terms, suppressed as part of ()(***) in (1.2), is a delicate task. For the fourth moment of the $\zeta$-function, Heath-Brown [34] and Motohashi [59] proved the asymptotic formula
@@
with $\delta<1/8$ and $\delta<1/3$ respectively, where $P_{2}$ is a polynomial of degree $4$. The resulting expressions of $P_{2}$ (see also Conrey [25]) were, unfortunately, quite complicated, making a complete determination of its coefficients daunting from their formulae.
In [59], Motohashi considers the following smoothed and shifted variant of the fourth moment:
@@
The shifts are initially taken with $\operatorname{Re}\alpha_{i},\,\operatorname{Re}\beta_{j}\gg 1$, and play a pivotal role in an intricate analytic continuation procedure. CFKRS later discovered that the many degenerate terms in Motohashi’s formula for (1.4) can be combined, in rather subtle ways, to yield an elegant, intrinsic description of the main terms. There are six main terms in the CFKRS formulation (see (5.2)), from which the polynomial $P_{2}$ can be computed in full and with ease by letting the shifts $\alpha_{i},\beta_{j}\to 0$. This is explained in [22, Section 5.1].
The recipe works equally well for the cubic moment of $\hbox{GL}(2)$ $L$-functions. Owing to differences in the families and their arithmetic, this case yields eight distinct main terms. In [24], the authors located only the diagonal term, though they noted the interest in identifying the others ([24, p. 1177]). We review these conjectures in Section 5, and examine the difficulties that hindered the identification of these terms in earlier approaches in Section 1.4.
@@ 1.3. Beyond the Petersson/Kuznetsov paradigm
Motohashi’s original work [59] is technical, and the source of “reciprocity” is far from apparent. A priori, there is no reason to expect the off-diagonal (or error) terms to be neatly expressible by moments of $L$-functions. Having a conceptual understanding of this phenomenon, without intermediate devices like Kloosterman sums, summation formulae, shifted convolutions, or the Petersson–Kuznetsov formulae, is therefore of significant interest.
This direction was perhaps first pursued by Bruggeman–Motohashi [12], whose approach to constructing the two distinct-looking moments in (1.2) involves a combination of spectral projections, Hecke operators, and a two-fold limiting process applied to an automorphic kernel. The choice of the kernel is by no means straightforward: it entails delicate smoothing and regularization, along with careful control of convergence, regularity, and interchange of limits. These ingredients, in turn, illuminate the integral transform through the theories of Whittaker–Kirillov models and Bessel functions for the group $\hbox{PGL}_{2}(\mathbb{R})$ (see [17]).
Michel–Venkatesh [56, Section 4], [57, Section 4.5], and Reznikov [69, p. 453] took a further step towards an account more intrinsic than [12]. They envisioned proving Motohashi’s formula within the structural framework of the strong Gelfand formation and via a “regularized” geodesic period for the product of two $\hbox{PGL}(2)$ Eisenstein series, formally given by
@@
which morally suggests a matching of the spectra $L^{2}([\mathrm{PGL_{2}}])$ and $L^{2}([\mathrm{GL_{1}}])$, resulting in Motohashi’s formula. However, numerous technical obstacles must be overcome to implement their strategy. In particular, the sources of various main (or degenerate) terms are not evident, and many integrals suffer from serious divergence issues due to the non-compactness of domains and the non-square-integrability of Eisenstein series. More refined regularization techniques than those developed in [57] are therefore necessary; see also [7] for other intricacies of this approach. Only very recently did Nelson [62] establish full rigour for the Michel–Venkatesh–Reznikov period integral strategy, through a range of highly impressive and sophisticated techniques.
Wu [71] gave another interpretation of Motohashi’s formula by generalizing [12] with a pre-trace formula on the Schwartz space of 22-by-22 matrices, which incorporates the Godement–Jacquet and Tate $L$-functions on the spectral and geometric sides of the formula. Further explications and analytic subtleties were addressed in [3], which draws extensively on deep results in the theory of special functions. The authors of [3] discuss the obstacles to obtaining an exact spectral inversion of Motohashi’s formula.
@@ 1.4. Our main results
In this article, we present a new proof of the spectral inversion of Motohashi’s identity (1.2) via the $\hbox{GL}(3)$ period integral method of [49]. Our approach offers fresh perspectives on the reciprocity phenomena and surmounts technical obstacles in earlier treatments. Furthermore, we elucidate the links between periods and the Moment Conjectures of [22].
In our previous work [49] (resp. [50]), we examine the (twisted) spectral first moment of $\hbox{GL}(3)\times\hbox{GL}(2)$ Rankin–Selberg $L$-functions with the $\hbox{GL}(3)$ automorphic form being a fixed Hecke–Maass cusp form. We uncover a higher-rank Motohashi-type identity that underpins Li’s celebrated convexity-breaking results [52].
This line of works belong to the recent framework Period Reciprocity, for which only a handful of other instances are known: [57, 9, 62, 63, 75, 76, 45, 55]. The tools involved lie largely outside the traditional realm of analytic number theory, and they streamline the intricate chains of analytic-arithmetic transformations prevalent in the literature. We refer the reader to [49, Sections 3-4], [50, Section 1] and [45] for more in-depth discussions of the implementations and technical features of this framework.
To state our main results, let $\alpha\in\mathfrak{a}_{\mathbb{C}}:=\{(\alpha_{1},\alpha_{2},\alpha_{3})\in\mathbb{C}^{3}:\alpha_{1}+\alpha_{2}+\alpha_{3}=0\}$, and $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ denote the following shifted cubic moment: [^1] The shifts are labelled with minus signs for automorphic reasons, as we will discuss in the main body of this article.
@@
where $(\phi_{j})_{j=1}^{\infty}$} is an orthogonal basis of even Hecke–Maass cusp forms of $\mathrm{SL}_{2}(\mathbb{Z})$ with respect to the Petersson inner product $\langle\,\,\cdot\,$$\cdot\,\rangle$ (see (2.9)), and satisfies $\Delta\phi_{j}=(1/4-\mu_{j}^{2})\phi_{j}$} with $\Delta:=-y^{2}(\partial_{x}^{2}+\partial_{y}^{2})$. Also, let $\Lambda(s)$ and $\Lambda(s,\,\phi_{j})$ be, respectively, the completed $\zeta$-function and completed $L$-function of $\phi_{j}$.
To simplify our residual calculations (as in [22]), we introduce:
@@
In practical applications of any spectral summation formula, it is essential to have an ample supply of admissible test functions. The following regularity assumption serves our purposes and is satisfied by those commonly used in the literature (cf. [49, Remark 5.27]).
@@
In the following, let $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\,\Gamma(s/2)$; $d^{*}\mathbf{y}:=(y_{0}y_{1}^{2})^{-1}\,dy_{0}\,dy_{1}$;  $e(x):=e^{2\pi ix}$; $H^{\flat}:(0,\infty)\rightarrow\mathbb{C}$ denote the inverse Kontorovich–Lebedev transform of $H$ (Definition 3.8); $d^{W}\mu:=|\Gamma_{\mathbb{R}}(\mu)|^{-2}\,\frac{d\mu}{4\pi i}$ be the spherical Whittaker–Plancherel measure for $\mathrm{PGL}_{2}(\mathbb{R})$; and $W_{-\alpha}(x,y_{1})$ be the Jacquet–Whittaker function for $\mathrm{PGL}_{3}(\mathbb{R})$ (see (3.6)).
@@ 1.4.1. The reciprocity formula
$\mathfrak{M}_{-\alpha}^{(3)}(s;H)\ +\ \mathcal{R}_{-\alpha}(s;H)\ +\ \mathcal{R}_{\alpha}(1-s;H)$
$\hskip 50.0pt\ =\ \sum_{i=1}^{3}\,\mathcal{M}_{-\alpha}^{i}(s;H)\,+\ \frac{1}{2}\ \int_{(1/2)}\zeta\left(2s-s_{0}\right)\ \prod_{i=1}^{3}\ \zeta(s_{0}+\alpha_{i})\left(\mathcal{F}_{\alpha}H\right)\left(s_{0},s\right)\ \frac{ds_{0}}{2\pi i},$ (1.7)
@@
$(\mathcal{F}_{\alpha}H)(s_{0},s)\ :=\ \sum_{\pm}\,\int_{0}^{\infty}\int_{0}^{\infty}\,H^{\flat}\Big(\frac{y_{1}}{\sqrt{1+y_{0}^{2}}}\Big)\int_{0}^{\infty}W_{-\alpha}(x,y_{1})e(\pm xy_{0})x^{s_{0}-1}\,d^{\times}x\,\frac{y_{0}^{2s-s_{0}}y_{1}^{s-\frac{1}{2}}}{(1+y_{0}^{2})^{\frac{s}{2}+\frac{1}{4}-s_{0}}}\,d^{*}\mathbf{y},$ (1.8)
$\mathcal{M}_{-\alpha}^{0}(s;H)\,$ $:=\,\prod_{1\leq i<j\leq 3}\,\zeta(2s-\alpha_{i}-\alpha_{j})\ \int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i})\ d^{W}\mu,$ (1.9)
$\mathcal{M}_{-\alpha}^{1}(s;H)\,$ $:=\,\sum_{i=1}^{3}\,\ \zeta(2s-\sum_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\,\alpha_{j})\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\,\zeta(1+\alpha_{i}-\alpha_{j})$
$\hskip 80.0pt\times\ \int_{(0)}\,H(\mu)\prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{i})\prod_{\begin{subarray}{c}j\neq i\end{subarray}}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{j})\,d^{W}\mu,$ (1.10)
$\mathcal{M}_{-\alpha}^{2}(s;H)\,$ $:=\,\ \sum_{i=1}^{3}\ \zeta\big(2-2s+\sum_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\alpha_{j}\big)\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\zeta(1-\alpha_{i}+\alpha_{j})$
$\hskip 80.0pt\,\times\,\int_{(0)}\,H(\mu)\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i})\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{j})\ d^{W}\mu,$ (1.11)
$\mathcal{M}_{-\alpha}^{3}(s;H)\,$ $:=\,\ \prod_{1\leq i<j\leq 3}\,\zeta(2-2s+\alpha_{i}+\alpha_{j})\int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\ \prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{i})\,d^{W}\mu,$ (1.12)
$\mathcal{R}_{\alpha}(s;H)\$:= $:=\ 2\ \sum_{i=1}^{3}\ H(s+\alpha_{i})\ \frac{\prod_{\begin{subarray}{c}1\leq j\leq 3\\ j\neq i\end{subarray}}\ \Lambda(1+\alpha_{i}-\alpha_{j})\Lambda(2s+\alpha_{i}+\alpha_{j})}{\Lambda\left(1+2s+2\alpha_{i}\right)}.$ (1.13)